Find the Fibonacci number when n= 10 (𝐹10), using recursive relation. Take 𝐹0=2 and 𝐹1 =5. Upload your solution.
Answer by ikleyn(52780)

.

Use the formula for the Fibonacci sequense

    F(n+2) = F(n) + F(n+1).


The second term is  F(2) = F(1) + F(0) = 5 + 2 = 7;


The third term is   F(3) = F(2) + F(1) = 7 + 5 = 12


and so on.


Continue these calculations until you will get the 10-th term.


This exercise is a good job for a student, who wants to develop his (or her) skills and knowledge.

Happy calculations (!)